Nitrogen and hydrogen combine at high temperature,in presence of a catalyst,to produce ammonia.N2(g)+3H2(g)=2NH3.Assume 0.100 mol of N2 and 0.329 mol of H2 are present initially. After complete reaction,how many mols of ammonia are produced?How many moles of H2/N2 remain?
Answer #1

mol of N2 reqruired for 0.329 mol --> 1/3 * 0.329 = 0.109666 mol of N2, which we do not have, so N2 is limiting

mol of N2 reacted = 0.1

mol of N2 left = 0.1-0.1 = 0

mol of H2 reacted = 0.1*33 = 0.3 mol

mol of H2 left = 0.329 - 0.3 = 0.029 mol of H2

mol of NH3 produced = 0.1*2 = 0.2 mol of NH3

so..

total mol = 0.029+0.2 = 0.229 mol finally